A student wrote 69 as the average of the data: 70,59,72,35,67,78,68,81,80,and 76. By how much was the answer incorrect?
Found 2 solutions by checkley75, tutorcecilia:
Answer by checkley75(3666)

70+59+72+35+67+78+68+81+80+76=552
average=552/10=55.2
incorrect by 69-55.2=13.8 points.

Answer by tutorcecilia(2144)

Find the average:
(70+59+72+35+67+78+68+81+80+76)/10 = 68.6
.
69-68.60 = 0.4 [The student was off by 0.4]